# TSA-WF: Exploring the Effectiveness of Time Series Analysis for Website Fingerprinting

### v1

@@ 3.1 Representing Traces as Time Series
In preliminary experiments we observed that, when algorithms based on time series similarity metrics (e.g., euclidean distance) attempt to match time series represented in this format, they tend to produce unrepresentative similarity scores. This is due to two main factors: a) the representation makes for abrupt jumps in the value of $y$ when switching between incoming and outgoing packets, and; b) if there are horizontal delays (e.g., due to packet re-transmission) between two traces, this difference is reflected in both the positive and negative components of the time series. Existing work using time series analysis for WF has not addressed these issues, leveraging this simple representation [30, 25].
@@ 3.2 Classical Time Series Matching for WF
The first of such techniques is euclidean distance, which simply measures the straight-line difference between corresponding points in two sequences, treating each point as a dimension in a multi-dimensional space. Euclidean distance is deemed simple, accurate, and effective but can be computationally intensive [23]. This technique has been used in the past to detect attacks in network traffic [44]. The second technique we consider is STUMPY [21], which aims to make time series comparison fast and efficient by computing an approximation of the euclidean distance between two time series using a matrix profile [42] along with further optimizations. STUMPY has been previously used for time series anomaly detection [12]. Finally, compression-based-distance (CBD) [19] is an indirect approach to comparing the distance between two time series. In CBD, time series are converted into symbols (e.g. with symbolic aggregate approximation [13]) that can be stored as a text file. Then, these files are compressed (e.g., .zip, .rar) and the difference in their sizes are compared. Compression-based metrics have been used before for anomalous web traffic analysis [33].
@@ 3.3 Challenges for WF via Time Series Analysis
The two issues identified above with (non-)normalized euclidean distance call for the consideration of an additional time series comparison metric which can complement euclidean-based distance calculations. To this end, when implementing TSA-WF (§4.2), we explore the use of CBD distance calculations which eschew the magnitude of $y$ values (mitigating the non-normalization issue), without homogenizing different traces (mitigating the normalization issue).
@@ 4 The Architecture of TSA-WF
In TSA-WF’s execution pipeline (shown in Fig. 7), the adversary first obtains prototypes, which are representative samples for each website in the monitored set. Then, during distance computation and training, these prototypes are compared with samples from the training set using time series distance metrics to train a classifier. Finally, the classifier is evaluated against unlabeled samples based on its ability to predict the monitored website and untangle the multi-tab trace. Next, we detail each of these phases.
@@ 4.2 Distance Computation and Training
The second phase of TSA-WF computes the distance between the (few) selected prototypes from each website (i.e., class) and all other trace samples in the training data. These distances will be later used to train a classifier assuming that the distances from a class’ prototypes to samples of its own class should be smaller that the distances obtained when comparing the prototype to samples which belong to other classes. TSA-WF combines the distances computed with euclidean (with and without weights), STUMPY, CBD, and DTW by representing trace similarity as a vector $n$ with five cells (one for each distance metric we consider).
Calculating distances. Algorithm 1 details the computation of the distance between a class prototype and a training sample. Function ComputeDist (line 4) aims to find the minimum distance between a prototype $s$ and sample $t$ (both stored as lists containing the trace representation). Specifically, ComputeDist finds the subsequence of $t$ that is most similar to $s$ (line 8) using each similarity function stored in $n$ (line 3) independently, then concatenates the final result. For every distance function in $n$ (line 6), we use a sliding window with size $len(s)$ and step size 1 (line 7) that moves across $t$ and computes the distance at each interval to produce a matrix of distances $d$ with shape $[len(t)-len(s),len(n)]$ (line 5). In the (unlikely) scenario that the prototype length exceeds the trace length, we reverse $t$ and $s$. Then, TSA-WF reduces this matrix into a vector with shape $[n]$ by returning the minimum value of each distance metric (line 9).
@@
6:    for $j=0\rightarrow len(n)$, compute_dist = n[j] do $\triangleright$ Metrics
@@
9:    return min($d$[:,$j$]) $\triangleright$ Minimum of each distance metric
@@ 5.1 Experimental Setup
Trace distance metrics. To compute the distance between websites’ prototypes and other traces contained in TSA-WF’s training/testing data, we use existing code for STUMPY, the numpy implementation of euclidean distance, the tslearn implementation for DTW, and our own implementation of CBD.
@@ 5.2 Parameter Tuning
When performing these experiments, we used STUMPY as our default distance metric (except when configuring weights for an actual implementation of euclidean distance) and a random forest classifier (except when testing alternative models to use in the final prediction step of TSA-WF). We experiment (and set) each of the aforementioned parameters in a cascade fashion.
@@
When evaluating TSA-WF using these variations of WED in lieu of STUMPY, we obtained an accuracy of 83.5% for linear weights, 84.8% for logarithmic weights, and 85.1% for reflected logarithmic weights. While WED does not excel as an independent distance metric when compared to STUMPY, its combination with other metrics can enhance TSA-WF’s accuracy (see §5.3). For the remainder of our experiments, we report WED using reflected logarithmic weights.
@@ 5.3 TSA-WF’s Effectiveness on Single-Tab Traces
Classification using individual distance metrics. Initially, we evaluate our time-series method by employing each distance metric independently. STUMPY exhibits the highest classification accuracy (87.9%) when compared with all other time series analysis methods (i.e., euclidean, WED, CBD, and DTW).
Classification using a combination of distance metrics. As discussed in §3.3, each similarity metric captures different aspects of website trace matching and confers benefits in certain respects while lacking in others. As anticipated, TSA-WF’s accuracy increased when the pipeline’s prediction model is supplied with the feature vectors obtained from multiple metrics combined. Indeed, combining STUMPY, euclidean, WED, CBD and DTW, results in an accuracy of 92.2% – offering an improvement of 4.3% over the use of STUMPY alone.
Benchmarking WF attacks on undefended traces. We also benchmarked a set of prominent WF attacks on the same dataset used to evaluate TSA-WF above. As shown in Table 4, TSA-WF’s classification pipeline (fueled by all distance metrics) outperformed Tik-Tok, DF, and k-FP, distancing itself from the accuracy achieved by k-FP by 2.4%.
@@ 5.4 TSA-WF’s Effectiveness on Multi-Tab Traces
Merged traces. In this experiment, we compare the performance of TSA-WF with existing work designed for the multi-tab setting in the open-world scenario. First, similarly to our observations in the single-tab setting, we can see in Table 4 that the standalone use of STUMPY in TSA-WF’s pipeline still reaps most of the benefits when compared to the combined usage of distance metrics. Specifically, we can observe that TSA-WF (STUMPY) achieves an accuracy of 50.5% for the 3-Tab scenario, while the combination of STUMPY, WED, euclidean, and CBD distances achieves an accuracy of 52.2%, only 1.7% better. However, we also see that, in the multi-tab setting, CBD helped improve TSA-WF’s accuracy, as opposed to the single-tab setting where its use provided no advantage.

### v2

@@ 3.1 Representing Traces as Time Series
In preliminary experiments we observed that, when algorithms based on time series similarity measures (e.g., euclidean distance) attempt to match time series represented in this format, they tend to produce unrepresentative similarity scores. This is due to two main factors: a) the representation makes for abrupt jumps in the value of $y$ when switching between incoming and outgoing packets, and; b) if there are horizontal delays (e.g., due to packet re-transmission) between two traces, this difference is reflected in both the positive and negative components of the time series. Existing work using time series analysis for WF has not addressed these issues, leveraging this simple representation [30, 25].
@@ 3.2 Classical Time Series Matching for WF
The first of such techniques is euclidean distance, which simply measures the straight-line difference between corresponding points in two sequences, treating each point as a dimension in a multi-dimensional space. Euclidean distance is deemed simple, accurate, and effective but can be computationally intensive [23]. This technique has been used in the past to detect attacks in network traffic [44]. The second technique we consider is STUMPY [21], which aims to make time series comparison fast and efficient by computing the euclidean distance between two time series using a matrix profile [42] along with further optimizations. STUMPY has been previously used for time series anomaly detection [12]. Finally, compression-based-distance (CBD) [19] is an indirect approach to comparing the distance between two time series. In CBD, time series are converted into symbols (e.g. with symbolic aggregate approximation [13]) that can be stored as a text file. Then, these files are compressed (e.g., .zip, .rar) and the difference in their sizes are compared. Compression-based measures have been used before for anomalous web traffic analysis [33].
@@ 3.3 Challenges for WF via Time Series Analysis
The two issues identified above with (non-)normalized euclidean distance call for the consideration of an additional time series comparison measure that can complement euclidean-based distance calculations. To this end, when implementing TSA-WF (§4.2), we explore the use of CBD distance calculations which eschew the magnitude of $y$ values (mitigating the non-normalization issue), without homogenizing different traces (mitigating the normalization issue).
@@ 4 The Architecture of TSA-WF
In TSA-WF’s execution pipeline (shown in Fig. 7), the adversary first obtains prototypes, which are representative samples for each website in the monitored set. Then, during distance computation and training, these prototypes are compared with samples from the training set using time series distance measures to train a classifier. Finally, the classifier is evaluated against unlabeled samples based on its ability to predict the monitored website and untangle the multi-tab trace. Next, we detail each of these phases.
@@ 4.2 Distance Computation and Training
The second phase of TSA-WF computes the distance between the (few) selected prototypes from each website (i.e., class) and all other trace samples in the training data. These distances will be later used to train a classifier assuming that the distances from a class’ prototypes to samples of its own class should be smaller that the distances obtained when comparing the prototype to samples which belong to other classes. TSA-WF combines the distances computed with euclidean (with and without weights), STUMPY, CBD, and DTW by representing trace similarity as a vector $n$ with five cells (one for each distance measure we consider).
Calculating distances. Algorithm 1 details the computation of the distance between a class prototype and a training sample. Function ComputeDist (line 4) aims to find the minimum distance between a prototype $s$ and sample $t$ (both stored as lists containing the trace representation). Specifically, ComputeDist finds the subsequence of $t$ that is most similar to $s$ (line 8) using each similarity function stored in $n$ (line 3) independently, then concatenates the final result. For every distance function in $n$ (line 6), we use a sliding window with size $len(s)$ and step size 1 (line 7) that moves across $t$ and computes the distance at each interval to produce a matrix of distances $d$ with shape $[len(t)-len(s),len(n)]$ (line 5). In the (unlikely) scenario that the prototype length exceeds the trace length, we reverse $t$ and $s$. Then, TSA-WF reduces this matrix into a vector with shape $[n]$ by returning the minimum value of each distance measure (line 9).
@@
6:    for $j=0\rightarrow len(n)$, compute_dist = n[j] do $\triangleright$ Measures
@@
9:    return min($d$[:,$j$]) $\triangleright$ Minimum of each distance measure
@@ 5.1 Experimental Setup
Trace distance measures. To compute the distance between websites’ prototypes and other traces contained in TSA-WF’s training/testing data, we use existing code for STUMPY, the numpy implementation of euclidean distance, the tslearn implementation for DTW, and our own implementation of CBD.
@@ 5.2 Parameter Tuning
When performing these experiments, we used STUMPY as our default distance measure (except when configuring weights for an actual implementation of euclidean distance) and a random forest classifier (except when testing alternative models to use in the final prediction step of TSA-WF). We experiment (and set) each of the aforementioned parameters in a cascade fashion.
@@
When evaluating TSA-WF using these variations of WED in lieu of STUMPY, we obtained an accuracy of 83.5% for linear weights, 84.8% for logarithmic weights, and 85.1% for reflected logarithmic weights. While WED does not excel as an independent distance measure when compared to STUMPY, its combination with other measures can enhance TSA-WF’s accuracy (see §5.3). For the remainder of our experiments, we report WED using reflected logarithmic weights.
@@ 5.3 TSA-WF’s Effectiveness on Single-Tab Traces
Classification using individual distance measures. Initially, we evaluate our time-series method by employing each distance measure independently. STUMPY exhibits the highest classification accuracy (87.9%) when compared with all other time series analysis methods (i.e., euclidean, WED, CBD, and DTW).
Classification using a combination of distance measures. As discussed in §3.3, each similarity measure captures different aspects of website trace matching and confers benefits in certain respects while lacking in others. As anticipated, TSA-WF’s accuracy increased when the pipeline’s prediction model is supplied with the feature vectors obtained from multiple measure combined. Indeed, combining STUMPY, euclidean, WED, CBD and DTW, results in an accuracy of 92.2% – offering an improvement of 4.3% over the use of STUMPY alone.
Benchmarking WF attacks on undefended traces. We also benchmarked a set of prominent WF attacks on the same dataset used to evaluate TSA-WF above. As shown in Table 4, TSA-WF’s classification pipeline (fueled by all distance measures) outperformed Tik-Tok, DF, and k-FP, distancing itself from the accuracy achieved by k-FP by 2.4%.
@@ 5.4 TSA-WF’s Effectiveness on Multi-Tab Traces
Merged traces. In this experiment, we compare the performance of TSA-WF with existing work designed for the multi-tab setting in the open-world scenario. First, similarly to our observations in the single-tab setting, we can see in Table 4 that the standalone use of STUMPY in TSA-WF’s pipeline still reaps most of the benefits when compared to the combined usage of distance measures. Specifically, we can observe that TSA-WF (STUMPY) achieves an accuracy of 50.5% for the 3-Tab scenario, while the combination of STUMPY, WED, euclidean, and CBD distances achieves an accuracy of 52.2%, only 1.7% better. However, we also see that, in the multi-tab setting, CBD helped improve TSA-WF’s accuracy, as opposed to the single-tab setting where its use provided no advantage.